# Large Language Models Synergize with Automated Machine Learning

### v1

@@ 6.1 Setup
To explore the limits of program synthesis for ML, in experiments, we consider the automatic code generation of the entire ML pipeline, given textual descriptions of the ML task. For an ML task, let the set of valid programs (as defined in Section 6.3) for the entire ML pipeline of the task be the desired sequences $\Omega$. Given the textual description $\mathbb{T}$, validation data $D^{val}$, and evaluation metrics $\mathbb{E}$, then the problem is to automatically generate the program $\mathbf{C}_{\theta}$ that maximizes the evaluation score based on $\mathbb{E}$ and $D^{val}$:
$\operatorname*{arg\,max}_{\theta}\mathbb{E}\,(\mathbf{C}_{\theta},D^{val})\quad\text{s.t.}\;\mathbf{C}_{\theta}\in\Omega$
In other words, at first, a precondition is that the program is generated and is free of major syntax errors. Then for all validly generated programs, the goal is to select the best program, by maximizing the performance score evaluated by the metrics. For other sequence generation methods, which are incompatible with autoML due to lack of external compatibility (see Section 4), we only evaluate their efficiency in generating valid programs (Section 6.2).
For our method, we further perform optimization (Section 6.4). We generate each candidate program $\mathbf{C}_{\theta}$ in an iterative approach through LLMs and autoML. Given an LLM $\lambda$, generation method $\Psi:\lambda\times\mathbb{T}\times\theta\rightarrow\mathbf{C}_{\theta}^{*}$, autoML method $\Theta$, training data $D^{tr}$, and training procedure $\mathbb{H}:\mathbf{C}_{\theta}^{*}\times D^{tr}\times\theta\rightarrow\mathbf{C}_{\theta}$, in the $i_{th}$ iteration, our method generates $\mathbf{C}_{i}$:
$\mathbf{C}_{i}=\mathbb{H}\,(\mathbf{C}_{\theta}^{*},D^{tr},\Theta(i))\text{, where}\;\mathbf{C}_{\theta}^{*}=\Psi\,(\lambda,\mathbb{T},\Theta(i))$
@@
Setup of Methods We use GPT-4, GPT-3.5, Achiam et al. (2023) and PaLM 2 Anil et al. (2023) as LLMs $\lambda$. In Algorithm 1, we incorporate the self-reflection mechanism in Shinn et al. (2023) in $\{\Phi^{j}\}$. We employ both random search and the hyperparameter optimization method BOHB Falkner et al. (2018) as the search strategy $\mathbb{H}$ (Appendix C.3). For ZC proxies, we utilize flops, params, naswot, and synflow (Appendix C.5). flops and params are two simple yet powerful ZC proxies Krishnakumar et al. (2022). While most ZC proxies are designed for CV tasks, Zhou et al. (2022b) and Javaheripi et al. (2022) suggest the suitability of naswot and synflow for NLP tasks. Throughout the experiments, for the same dataset, we provide the same task description $\mathbb{T}$ for each generation method (Appendix E).
@@ 6.2 Generation of valid programs
As outlined in the task of repeated generations (Section  3), we compare each sequence generation method across the datasets by repeatedly attempting to generate valid ML programs. This process is capped at a maximum of 100 attempts. If a method generates a valid program within these 100 attempts, the method is reset and tasked with generating another valid program until all 100 attempts are exhausted.
Among the techniques, Reflexion and Contextual Modular Generation incorporate a self-reflection mechanism, as outlined by Shinn (2023). This mechanism retains memories of previous trials, including feedback from testing results of earlier iterations (see Figure 3 and Algorithm 1). Such feedback is used as an input sequence to provide insights for the generation process. However, we observed that these two methods often fall into repetitive cycles involving the same code-errors-feedback loop. To address this issue, we force a memory reset if the method fails to produce a valid program after 10 attempts, thus mitigating the effects of poor initiations.
Table 1: Mean number of attempts per valid program with GPT-4. (1) CoT, Zero-shot, and Reflexion refer to Wei et al. (2022), Kojima et al. (2022), and Shinn et al. (2023) respectively. (2) Self-Revision: the generation based on the replacement of the contextual modular evaluations in Algorithm 1 with the self-revision mechanism in Le et al. (2023). (3) for Contextual Modular Generation, w/o tests: Algorithm 1 by removing the unit tests (Figure 7 and Algorithm 3), w/o reflection: Algorithm 1 by removing the self-reflection mechanism in $\{\Phi^{j}\}$, standard: Algorithm 1. (4) the datasets are divided into three groups in the order of tasks for tabular data, CV tasks, and NLP tasks. Details of each dataset are available in Appendix D. (5) -: the method is unable to generate any valid program for the dataset within 100 attempts.
@@
We also conduct experiments with GPT-3.5 and PaLM 2. For both LLMs, Text-to-ML outperforms other methods in 11 out of the 12 datasets (Appendix F). Overall, GPT-4 demonstrates the highest success rate among the LLMs.
@@ 6.3 Unit tests and false positives
In a strict sense, errors in ML programs are nearly impossible to eliminate compared to traditional coding problems. Traditionally, software systems are constructed deductively by explicitly programming the rules that govern system behavior. In contrast, ML rules are inferred inductively from the data Braiek & Khomh (2020). For example, the program may be error-free for one set of data points but buggy for another Kim et al. (2019); Braiek & Khomh (2020); Wang et al. (2021). Moreover, many ML programs rely on high-level libraries like PyTorch and TensorFlow, and third-party libraries such as the Intel Math Kernel Library, which often lack comprehensive specifications or even detailed source code documentation Braiek & Khomh (2020).
Nevertheless, as in the practices of many human programmers, ML programs can still be free of major errors and function effectively in most daily scenarios. In our study, we define valid programs as those that satisfy two criteria: (1) the program executes over the datasets in the task without interruptions from errors detectable by the execution environment (the Python interpreter), and (2) for programs that meet the first requirement, we manually ensure that the programs form a complete pipeline for the specified ML task.
Similar to traditional coding problems Li et al. (2022), an ML program can be a false positive, appearing valid in unit tests but actually failing to meet the criteria mentioned above. In our study, each program is divided into modules, each verified by unit tests before assembly. However, modules verified as valid individually may not necessarily remain valid when combined. Therefore, the entire assembled program undergoes an additional simple test by the execution method to automatically verify the first criterion.
In the experiment in Section 6.1, for our method (Contextual Modular Generation - standard Table 2), we examine the false positive rates of the generated programs both before and after this additional testing phase.
Table 2: Unit tests efficiency and false positive rates. (1) Traditional: tasks based on Traditional ML algorithms such as random forest and SVM. (2) FP rate w/o execution:False positive rates of pipeline code post-verification but prior to testing by the execution environment (3) FP rate w/o execution’: false positive rates of pipeline code following both verification and testing by the execution environment.
Modality Pipelines verified by unit tests FP rate w/o execution FP rate w execution
@@
As Table 2 shows, the false positive rates for deep learning tasks are significantly higher than those of traditional algorithms, but they are still within a reasonable range. The false positive rates of the programs after meeting the first criterion are noticeably low, which suggests the reliability of automatically created ML programs
Notably, we do not insist on the compatibility of generated modules with any pre-written module, which is more lenient towards the other methods, particularly those without testing, such as CoT and Zero-Shot. Achieving such external compatibility (Section 4), is infeasible for methods that do not include testing or rely on traditional testing techniques. However, in our method, the generated modules are also tested for compatibility, which is crucial for performing consistent training procedures and for a uniform evaluation of these modules.
@@ 6.4 Improvements from optimization
Our method ensures that generated programs are compatible with a consistent evaluation procedure. Leveraging this unique advantage, Text-to-ML automatically searches the programs generated by Contexual Modular Generation against metrics specified in the textual task descriptions. For each dataset, we generate codes of up to 20 candidate methods, each for data preparation and modeling and at least 1 code for post processing, forming up to 400 module combinations due to the internal connectivity (Section 4) provided by the testing technique (the network in Figure 7).
From these modules, we first randomly sample and evaluate 200 configurations of code combinations (pathways in the network in Figure 7) and their hyperparameters. We employ two optimization strategies: random search and BOHB (Falkner et al. (2018)) (Appendix C.3). For both strategies, we examine the identified program at the cost of 25 and 100 full evaluations (i.e., the cost equivalent to the cost of fully training and testing 25 and 100 models). In the search, ZC proxies and BOHB prematurely terminate the evaluations for less promising configurations (Appendix C.3). Then, we rank the best program identified during the search among the 200 sampled configurations.
For deep learning tasks, we train each model on 1 of 4 NVIDIA A100 GPUs. For sampling the configurations and random search, we train each model until it meets the early stopping criteria with a $patience$ of $3$ and a $minimum\;delta$ of $0.0$ Prechelt (2002).
The modules generated by traditional sequence generation methods are incompatible with autoML; however, these methods do allow for the configuration to be specified by LLMs. As a simple baseline, we also evaluate a single ML program chosen by the LLM without any further optimization. We utilize the prompt select the best configuration for the machine learning task (No Search in Table 3).
Table 3: Mean rank of the identified program. (1) Traditional: tasks based on Traditional ML algorithms such as random forest and SVM. (2)Rows with white and grey backgrounds represent the search with cost@25 and cost@100 respectively.
@@ 7 Discussion
By combining LLMs and autoML, we advance the frontier of program synthesis for ML. For the 12 tasks across tabular data, CV, and NLP, Text-to-ML, when provided with textual descriptions of tasks, autonomously generates optimized programs for the entire ML pipeline in a seamless, fully autonomous operation (Figure 1). Our study reveals the possibility of resolution of ML tasks as multi-step seq2seq tasks. Future research may explore such multi-step seq2seq tasks as autonomous agents. An interesting future exploration is the incorporation of additional text-to-text or image-to-text methods to automatically generate textual task descriptions, further automating the process.
From the perspective of autoML, our approach shifts the emphasis from enhancing the performance of optimized configurations on an algorithmic level to augmenting the automation of the ML workflow, particularly for everyday scenarios with limited resources. Text-to-ML is a potential tool for democratizing ML by significantly reducing the required efforts, expertise, and computational resources. Users simply provide a textual description of the task, and from there, the process is fully automatic, making ML in practice accessible to a broader audience.
@@
Limitations While the complexity of Contextual Modular Generation increases linearly over the problem length (Theorem 3.5), implementing this framework demands considerable effort (Appendix C.1). Furthermore, our approach still struggles at highly challenging ML tasks, particularly those that involve various forms of input and output data (Appendix D). Whether addressing these tasks requires a more efficient implementation of the framework or a transition to a different framework remains an open problem.
@@ 8 Conclusion
This study extends the scope of program synthesis for ML. By combining LLMs and AutoML, we achieve automatic generation of optimized programs of the entire ML pipeline over 12 tasks across traditional algorithms, CV, and NLP. Over 10 out of the 12 tasks, our generation framework Contextual Modular Generation outperforms existing methods, particularly for deep learning tasks and tasks with complex pipelines.
The testing technique Constrained Generative Unit Testing is instrumental for the implementation of the framework (Contextual Modular Generation w/o tests in Table 1). As a distinctive advantage of Text-to-ML, the compatibility with pre-written modules, including evaluation procedures and autoML methods, enables significant increases in the performance of the programs via optimization. In addition, we demonstrate the effectiveness of ZC proxies for pre-trained models in both CV and NLP.

### v2

@@ 6.1 Setup
To explore the limits of program synthesis for ML, in experiments, we consider the synthesis of the entire ML program, given textual descriptions of the ML task. For an ML task, let the set of valid programs (as defined in Section 6.3) for the entire ML program of the task be the desired sequences $\mathbf{\Omega}$. Given the textual description $\mathbf{T}$, testing data $D^{tes}$, and evaluation metrics $\mathbf{E}$, then the problem is to automatically generate the program $\mathbf{C}_{\theta}$ that maximizes the evaluation score based on $\mathbf{E}$ and $D^{val}$:
$\operatorname*{arg\,max}_{\theta}\mathbf{E}\,(\mathbf{C}_{\theta},D^{tes})\quad\text{s.t.}\;\mathbf{C}_{\theta}\in\mathbf{\Omega}$
In other words, at first, a precondition is that the generated programs are free of major syntax errors. Then among these programs, select the best program, by maximizing the performance score evaluated by the metrics. For other sequence generation methods, which are incompatible with autoML due to the lack of external compatibility (see Section 4), we only evaluate their efficiency in code generation (Section 6.2).
For our method, we further perform code optimization (Section 6.4). We generate each candidate program $\mathbf{C}_{\theta}$ in an iterative approach through LLMs and autoML. Given an LLM $f_{\lambda}$, generation method $\Psi:f_{\lambda}\times\mathbf{T}\times\theta\rightarrow\mathbf{C}_{\theta}^{*}$, autoML method $\Theta$, training data $D^{tr}$, and training procedure $\mathbf{H}:\mathbf{C}_{\theta}^{*}\times D^{tr}\times\theta\rightarrow\mathbf{C}_{\theta}$, in the $i_{th}$ iteration, our method generates $\mathbf{C}_{i}$:
$\mathbf{C}_{i}=\mathbf{H}\,(\mathbf{C}_{\theta}^{*},D^{tr},\Theta(i))\text{, where}\;\mathbf{C}_{\theta}^{*}=\Psi\,(f_{\lambda},\mathbf{T},\Theta(i))$
@@
Setup of methods We use GPT-4, GPT-3.5, (Achiam et al., 2023) and PaLM 2 (Anil et al., 2023) as LLMs $f_{\lambda}$. In Algorithm 1, we incorporate the self-reflection mechanism in Shinn et al. (2023) in $\{\Phi^{j}\}$ (see Appendix A.1). We employ both random search and the hyperparameter optimization method BOHB (Falkner et al., 2018) as the algorithms for code optimization (see Appendix C.3). For ZC proxies, we utilize flops, params, naswot, and synflow (Appendix C.5). flops and params are two simple yet powerful ZC proxies (Krishnakumar et al., 2022). While most ZC proxies are designed for CV tasks, Zhou et al. (2022b) and Javaheripi et al. (2022) suggest the suitability of naswot and synflow for NLP tasks. Throughout the experiments, for the same dataset, we provide the same task description $\mathbf{T}$ for each generation method (Appendix E).
@@ 6.2 Generation of valid programs
As outlined in the task of repeated generations (Section 3), we compare each sequence generation method across the datasets by repeatedly attempting to generate valid ML programs. This process is capped at a maximum of 100 attempts. If a method generates a valid program within these 100 attempts, the method is reset and tasked with generating another valid program until all 100 attempts are exhausted. For all methods, particularly Contextual Modular Generation, each sequence generation process (each time line 8 in Algorithm 1 is executed) counts as an attempt.
In Table 1, among the techniques, Reflexion and Contextual Modular Generation incorporate a self-reflection mechanism, as outlined by Shinn (2023). This mechanism retains memories of previous trials, including feedback from the testing results of earlier iterations (see Figure 3 and Algorithm 1). Such feedback is used as an input sequence to provide insights for the generation process. However, we observe that these two methods often fall into repetitive cycles involving the same code-errors-feedback loop. To address this issue, we force a memory reset if the method fails to produce a valid program after 10 attempts, thus mitigating the effects of poor initiations. For CoT, we set the number of input-output examples as 3 for GPT-4 and GPT-3.5 (Table 6) and 2 for PaLM 2 (Table 7).
Table 1: Mean number of attempts per valid program with GPT-4. (1) CoT, Zero-shot, and Reflexion refer to Wei et al. (2022), Kojima et al. (2022), and Shinn et al. (2023) respectively. (2) Self-Revision: the generation based on the replacement of the contextual modular evaluations in Algorithm 1 with the self-revision mechanism in Le et al. (2023). (3) for Contextual Modular Generation, w/o tests: Algorithm 1 by removing the unit tests (Figure 8 and Algorithm 3), w/o reflection: Algorithm 1 by removing the self-reflection mechanism in $\{\Phi^{j}\}$, standard: Algorithm 1. (4) the datasets are divided into three groups in the order of tasks for traditional algorithms, CV, and NLP. Details of each dataset are available in Appendix D. (5) -: the method is unable to generate any valid program for the dataset within 100 attempts.
@@
We also conduct experiments with GPT-3.5 and PaLM 2 (Appendix F). For both LLMs, Text-to-ML outperforms other methods in 11 out of the 12 datasets. Overall, GPT-4 demonstrates the highest success rate among the LLMs.
@@ 6.3 Unit tests and false positives
In a strict sense, errors in ML programs are nearly impossible to eliminate compared to traditional coding problems. Traditionally, software systems are constructed deductively by explicitly programming the rules that govern system behavior. In contrast, ML rules are inferred inductively from the data (Braiek & Khomh, 2020). For example, the program may be error-free for one set of data points but buggy for another (Kim et al., 2019; Braiek & Khomh, 2020; Wang et al., 2021b). Moreover, many ML programs rely on high-level libraries like PyTorch and TensorFlow, and third-party libraries such as the Intel Math Kernel Library, which often lack comprehensive specifications or even detailed source code documentation (Braiek & Khomh, 2020).
Nevertheless, as in the practices of many human programmers, ML programs can still be free of major errors and function effectively in most daily scenarios. In our study, we define valid programs as those that satisfy two criteria: (1) the program executes over the datasets in the task without interruptions from errors detectable by the execution environment (the Python interpreter), and (2) for programs that meet the first requirement, we manually ensure that the programs form a complete program for the specified ML task. Specifically, we manually check whether the generated program includes a complete ML pipeline, including data preparation, modeling, and post processing, as specified by the LLM in Search Space Generation in Algorithm 3 (see Appendix C.4).
Similar to traditional coding problems (Li et al., 2022), an ML program can be a false positive, appearing valid in unit tests but actually failing to meet the criteria mentioned above. In our study, each program is divided into modules, each verified by unit tests before assembly. However, modules verified as valid individually may not necessarily remain valid when combined. Therefore, the entire assembled program undergoes an additional simple test by the execution method to automatically verify the first criterion.
In the experiment in Section 6.2, for our method (Contextual Modular Generation - standard Table 1), we examine the false positive (FP) rates of the generated programs both before and after this additional testing step.
Table 2: Unit tests efficiency and false positive rates. (1) Traditional: tasks based on traditional ML algorithms such as random forest and SVM. (2) FP rate w/o execution: false positive rates of the programs post-verification but prior to testing by the execution environment (3) FP rate w/o execution: false positive rates of the programs following both verification and testing by the execution environment.
Modality ML programs verified by unit tests FP rate w/o execution FP rate w execution
@@
As Table 2 shows, the false positive rates for deep learning tasks are significantly higher than those of traditional algorithms, but they are still within a reasonable range. The false positive rates of the programs after meeting the first criterion are noticeably low, which suggests the reliability of ML programs that can be created automatically (by employing the execution step as an additional testing step).
Notably, for the experiments in Section 6.2, we do not insist on the compatibility of generated modules with any pre-written module, which is more lenient towards the other methods, particularly those without testing, such as CoT and Zero-Shot. Achieving such external compatibility (Section 4), is infeasible for methods that do not include testing or rely on traditional testing techniques. However, in our method, the generated modules are also tested for compatibility, which is crucial for performing consistent training and testing procedures and for a uniform evaluation of these modules.
@@ 6.4 Improvements from optimization
Our method ensures that generated programs are compatible with a consistent evaluation procedure. Leveraging this unique advantage, Text-to-ML automatically searches the programs generated by Contextual Modular Generation against metrics specified in the textual task descriptions. For each dataset, with GPT-4, we generate codes of up to 20 candidate algorithms, each for data preparation and modeling and at least 1 code for post processing, forming up to 400 module combinations due to the internal connectivity (Section 4) provided by the testing technique (the network in Figure 8).
From these modules, we first randomly sample and evaluate 200 configurations of module combinations (pathways in the network in Figure 8) and their hyperparameters. We employ two optimization strategies: random search and BOHB (Falkner et al., 2018) (Appendix C.3). For both strategies, we examine the identified program at the cost of 25 and 100 full evaluations (i.e., the cost equivalent to the cost of fully training and testing 25 and 100 models). In the search, ZC proxies and BOHB prematurely terminate the evaluations for less promising configurations (Appendix C.3). Then, we rank the best program identified during the search among the 200 sampled configurations.
For deep learning tasks, we train each model on 1 of 4 NVIDIA A100 GPUs. For sampling the configurations and random search, we train each model until it meets the early stopping criteria with a $patience$ of $3$ and a $minimum\;delta$ of $0.0$ (Prechelt, 2002).
The programs generated by traditional sequence generation methods are incompatible with autoML; however, these methods do allow for the program to be specified by LLMs. As a simple baseline, we also evaluate the one ML program chosen by the LLM without any further optimization (No Search in Table 3). We utilize the prompt select the best configuration for the machine learning task.
Table 3: Mean rank of the identified ML program in 10 trials. (1) Traditional: tasks based on traditional ML algorithms such as random forest and SVM. (2)Rows with white and grey backgrounds represent the search with cost@25 and cost@100 respectively.
@@ 7 Discussion
7.1 Synergy between LLMs and autoML
A major goal of autoML is to democratize ML by reducing the amount of effort required by human experts (Thornton et al., 2013; Wang et al., 2021a; Xin et al., 2021; Van der Blom et al., 2021). AutoML has achieved substantial progress in recent years on an algorithm level (Zoph & Le, 2016; Liu et al., 2018). However, in practice, autoML methods are gradually becoming a tool for experts to potentially boost the performance of the workflow (Tornede et al., 2023; Van der Blom et al., 2021), due to the following three issues: (1) most autoML methods only select the solutions at the algorithm level. However, the programming required for implementing the algorithms can demand substantial human effort and expertise. (2) for deep learning, many autoML methods select model architectures, such as the number of convolutional layers, from scratch and train the models anew, which is prohibitively expensive for regular users in practical scenarios under low resource constraints (Öztürk et al., 2022). (3) many autoML methods do not offer a comprehensive and unified solution applicable across task modalities (e.g., traditional algorithms, CV and NLP).
In our study, autoML numerically evaluates and selects ML programs generated by LLMs. Conversely, LLMs offer a solution to all three issues of autoML. As demonstrated in Figure 1, Section 6, and Figure 9, LLMs can generate the programs for the entire ML workflow across task modalities, while directly utilizing pretrained models for deep learning. Without the code generation from LLMs, there would be no code for autoML to select and optimize in the first place.
The objectives of AutoML and program synthesis for ML are closely aligned, with significant overlap in their purposes and methods (see Section 2.3). Nonetheless, strictly speaking, the goal of our study is on program synthesis for ML, rather than AutoML per se. This distinction arises because most autoML methods do not address ML tasks at the code level. In this context, AutoML is employed as a step in our method (Figure 3).
For the goal of synthesizing ML programs, LLMs synergize with autoML: LLMs bridge the gap between algorithm-centered autoML and autoML in practice, and autoML addresses the complex numerical evaluations of the programs generated by LLMs. In this synergy, LLMs contribute extensive coding knowledge embedded in their weights, facilitating code generation, while AutoML brings robust mathematical rigor from numerical optimization, enabling code optimization.
7.2 Sequence generation and autonomous agent
From a sequence generation perspective, our work explores the resolution of ML problems as a multi-step sequence to sequence (seq2seq) task (Sutskever et al., 2014). The input sequence is the textual task description. The output sequence is the corresponding optimized program for the entire ML program. Intermediary steps include sub-seq2seq tasks based on LLMs and optimization via autoML.
@@ 7.2 Sequence generation and autonomous agent
Our method, Text-to-ML, seamlessly integrates all intermediary steps into a single fully autonomous process (Figure 3 and Algorithm 2). Future research may explore such multi-step seq2seq tasks as autonomous agents. Another interesting future exploration is the incorporation of additional text-to-text or image-to-text methods to automatically generate textual task descriptions, further automating the process.
7.3 Limitations
@@
While the complexity of Contextual Modular Generation could increase only linearly over the problem length (Theorem 3.8), implementing this framework demands considerable effort (Appendix C.1). Furthermore, our approach still struggles at highly challenging ML tasks, particularly those that involve highly complex forms of input and output data (Appendix D). Whether addressing these tasks requires a more efficient implementation of the framework or a transition to a different framework remains an open problem.
In particular, challenging tasks on the Kaggle platform are typically addressed through the collective efforts of thousands of participants. The participants can freely exchange ideas and solutions, greatly enhancing the efficiency of developing high-performance solutions. Additionally, the highest-performing publicly shared solution often becomes the default baseline, as many participants replicate it in their solutions. These practices make a meaningful comparison between the solutions from human participants and automated solutions infeasible. In practice, most ML tasks are tackled by only one or a few human programmers.
Regardless, our work is not intended to outperform the collaborative efforts of thousands of competition participants with automatic solutions. Instead, our focus is on reducing the amount of effort and knowledge required for writing ML programs in everyday scenarios. Qualitatively, our method achieves an unprecedented degree of automation for ML tasks at the code level. Additionally, our work demonstrates significant performance improvements in the generated ML programs from the automatic optimization process.
@@ 8 Conclusion
This study extends the scope of program synthesis for ML. By combining LLMs and AutoML, we achieve the automatic generation and optimization of the complete ML programs over 12 tasks across traditional algorithms, CV, and NLP. Experimentally, over 10 out of the 12 tasks, our generation framework Contextual Modular Generation outperforms existing methods, particularly for deep learning tasks and tasks outside the training corpus of the LLMs. Theoretically, Contextual Modular Generation offers a linear best-case complexity class for generation difficulty over sequence length.
The testing technique Constrained Generative Unit Testing is instrumental for the implementation of the framework. As a distinctive advantage of Text-to-ML, the compatibility with pre-written modules, including a consistent numerical evaluation procedure, enables significant increases in the performance of the programs via numerical optimization from autoML. In addition, we demonstrate the effectiveness of ZC proxies for pretrained models in both CV and NLP.